If a price of corn is $3.00 a bushel, 5,000 bushels would be demanded. If the price rises to $4.00 a bushel, 4,000 bushels would be demanded
a. What is the (arc) price elasticity of demand?
b. Based on this answer, if the price of corn rose to $5.00 a bushel, what would be the demand for corn?
c. If the price of corn decreased from $4.00 to $3.00 a bushel, what would be the change in total revenue for sellers of corn?
d. If the price of corn increased from $4.00 to $5.00 a bushel, what would be the change in total revenue for sellers of corn?
1
Expert's answer
2020-03-31T09:00:00-0400

a. The (arc) price elasticity of demand is:

"Ed = (4,000 - 5,000)\/(4 - 3)\u00d7(4 + 3)\/(4,000 + 5,000) = = -7\/9 = -0.78,"

so the demand is inelastic.

b. Based on this answer, if the price of corn rose to $5.00 a bushel, what then the demand for corn would increase in: "(5\/3 - 1)\u00d7(-7\/9) = -0.52" and become near 2400 bushels.

c. If the price of corn decreased from $4.00 to $3.00 a bushel, the new quantity would be "((-1\/4)\u00d7(-7\/9) + 1)\u00d74,000 = 4,778," so the change in total revenue for sellers of corn would be "4\u00d74,000 - 3\u00d74,778 = 1,666."

d. If the price of corn increased from $4.00 to $5.00 a bushel, the change in total revenue for sellers of corn would be:

"4\u00d74,000 - 5\u00d72,400 = -4,000."